how do i do (1,-1) and (4-,4) with the slope intercept form?
Answer by stanbon(75887)

how do i do (1,-1) and (4,-4) with the slope intercept form?
----
slope = (-4--1)/(4-1) = -3/3 = -1
======================================
Cheers,
Stan H.
==========